# and Refinement Evaluation of Multimodal Models

### v1

@@ 1 Introduction
Vision language models (VLMs), such as LLaVA [32], GPT-4V [39], and Gemini [46], have demonstrated remarkable progress across various tasks [54, 30, 9] by integrating large language models (LLMs) [48, 20] with visual encoders [12, 42]. However, VLMs can sometimes produce undesirable outputs, possibly due to omitting important details in images or misunderstanding the instructions, which prompts the need for the feedback-refining capability beyond the normal instruction-following ability. This capability enables VLMs to spontaneously refine their responses based on user feedback, as depicted in Fig. 1, enhancing the efficiency and smoothness of interactions between users and visual assistants.
@@
In doing so, we build FIRE, a dataset for Feedback Integration and Refinement Evaluation of VLMs. FIRE is composed of 1.1M high-quality multi-turn feedback-refinement conversations, derived from $27$ source datasets across a wide range of tasks, such as visual question answering [15], image captioning [7], OCR reasoning [38, 43], document understanding [17], math reasoning [34], and chart analysis [36]. To scale up the data collection, FIRE is collected in two stages. In the first stage, we randomly sample $\sim$100K image-instruction-response triplets from data sources. We use each triplet to instruct GPT-4V to simulate a dialogue between a student and a teacher: the student answers the question and the teacher provides feedback to help the student improve its answer. We filter out generated low-quality conversations, such as those with too many turns or no improvement, rendering 100K high-quality feedback-refinement conversations, named FIRE-100K. In the second stage, we fine-tune two LLaVA-NeXT [31] models on FIRE-100K: one is trained as a student to refine its answer with the feedback, and the other is trained as a teacher to generate feedback for the student’s answer. We simulate dialogues between the student and the teacher models using $\sim$1M data points from the sources, rending a split named FIRE-1M. In this case, the full FIRE dataset consists of 1.1M feedback-refinement conversations in two splits FIRE-100K and FIRE-1M.
@@
To comprehensively evaluate the feedback-refining capability of VLMs, we build FIRE-Bench that has 11K feedback-refinement conversations derived from 16 source datasets, including 8 seen datasets (test splits) from FIRE-100K and FIRE-1M, as well as 8 new datasets from recently-proposed popular multimodal benchmarks. Using FIRE-Bench, we design two evaluation settings: fixed dialogues and free dialogues. In fixed dialogues, we compare the model’s refined response with ground truth in the generated conversations in FIRE-Bench, given a fixed dialogue history. In free dialogues, we let the model freely interact with a teacher model about instructions in FIRE-Bench, and test how fast & how much the model can improve its answers based on the feedback provided by the teacher model.
@@
We develop FIRE-LLaVA by fine-tuning LLaVA-NeXT [31] on FIRE-100K and FIRE-1M. The evaluation results on FIRE-Bench shows that FIRE-LLaVA exhibits significant improvements based on feedback in conversations, exceeding the original LLaVA-NeXT model method by $50\%$. These results underscore the significance of FIRE-100K and FIRE-1M in enhancing feedback integration, while FIRE-Bench provides an evaluation platform to analyze refinements. We expect that FIRE could motivate future exploration of the feedback-refining capability of VLMs.
@@
In summary, our contributions are three-fold. (1) We introduce FIRE, a dataset containing 1.1M feedback-refinement conversations across a wide range of tasks, where 100K data is generated by GPT-4V and 1M data is freely generated by simulating dialogues between student and teacher models. (2) We introduce the FIRE-Bench benchmark, composed of 11K conversations and a teacher model, providing comprehensive evaluations for the feedback-refining capability in two settings: fixed dialogues and free dialogues. (3) We develop FIRE-LLaVA, an advanced VLM that could improve its responses based on feedback, making efficient interaction between users and VLMs.
@@ 2.1 Vision Language Models
Building open-source VLMs to compete with closed-source models like GPT-4V [39] and Gemini [46] is a hot research topic. BLIP [24, 23] and Flamingo [1] are pioneering models that combine LLMs with visual encoders to enhance cross-modal understanding and reasoning abilities. LLaVA [32], InstructBLIP [11], and MiniGPT4 [54] develop the instruction tuning ability of VLMs by introducing a large number of instruction-response pairs. Along this way, some work focuses on the visual grounding ability of VLMs, such as Kosmos-2 [41], MINI-GPTv2 [5], and Qwen-VL [3], improving the region understanding for VLMs. InternVL [9] and mini-Gemini [26] develop powerful visual encoders for high-resolution images, and CuMo adopts a mixture-of-experts (MOE) architecture to manage diverse data better. Compared with existing VLMs, our FIRE-LLaVA has a more powerful feedback-refining capability across diverse tasks, which can spontaneously refine responses based on user feedback, leading to efficient and smooth interaction with users.
@@ 2.3 Feedback Learning in Multimodal Models
Learning from feedback is a promising research direction, playing an important role in human-robot interaction [27, 10]. Existing feedback learning methods can be roughly divided into two categories: planned feedback learning and impromptu feedback learning. Planned feedback learning updates models based on user feedback, and thus can generalize to new data but cannot provide refined responses immediately. CLOVA [14] and Clarify [21] are representative methods that automatically collect data to learn new knowledge. LLaVA-RLHF [45] collects human preference and trains VLMs via reinforcement learning. Impromptu feedback learning can immediately refine responses but have less generalization since they usually do not update models, which is widely studied in LLMs [2, 25, 47]. Liao et al. [28] use VLMs themselves as verifiers that produce feedback to correct recognition results. DRESS [8] generates helpfulness, honesty, and harmlessness responses via impromptu feedback learning. Different from DRESS, we improve the correctness and details of responses via impromptu feedback learning across diverse tasks.
@@ 3.1 Task Definition
Data Source. To enhance the diversity and comprehensiveness of our dataset, we compile more than 1.1M image-instruction-response triples from $27$ source datasets, being used to generate FIRE-100K, FIRE-1M, and FIRE-Bench, as shown in Fig. 2. These datasets cover tasks including visual question answering, image captioning, complex reasoning, OCR, chart/table/document analysis, math problems, science question answering etc.
@@ 4.2 Teacher Model
LLaVA-NeXT-8B 65.9 79.0 52.0 69.8 77.3 78.5 74.4 1546 331.4 44.9
FIRE-LLaVA 64.5 80.9 54.3 61.0 76.8 73.4 79.3 1548 340.5 38.3
@@ 5.5 Ablation Studies
In Fig. 7, we present the performance curve in FIRE-Bench concerning the number of turns in dialogues, evaluating LLaVA-NeXT, FIRE100K-LLaVA, and FIRE-LLaVA. We report the percentage of correctly answered samples (those scores greater than 8) after each turn. As the number of turns increases, the percentage of correctly answered samples rises across all three models. For FIRE-LLaVA, $49\%$ of the data is correctly answered in the first turn, and this increases to $69\%$ after five turns, with $20\%$ of the samples being corrected based on feedback. Compared to the LLaVA-NeXT model, FIRE-LLaVA shows an additional $7\%$ improvement (from $49\%\to 69\%$ vs. $46\%\to 59\%$), highlighting the effectiveness of FIRE-LLaVA when trained on FIRE.
@@ Appendix A Human Verification on FIRE
Figure A1: The screenshot of the interface for the human verfication.
Appendix B Additional Experimental Results
B.1 Error bar
@@
B.2 More VLMs
@@
B.3 More Visualization
We provide more cases tested on the FIRE-Bench to show the feedback-refining ability, as shown in Figs. A2, A3, A4, A5, A6, A7, A8 and A9. We observe that FIRE-LLaVA trained on FIRE has good feedback-refining ability on different tasks, such as visual perception, visual question answering, image captioning, OCR reasoning, document understanding, math reasoning, science question answering, and chart analysis.
@@ B.3 More Visualization
Appendix C Prompt
C.1 Prompt for GPT-4V
@@
C.2 Prompt for Student and Teacher models
@@
Appendix D Potential Negative Societal Impacts

### v2

@@ 1 Introduction
Vision language models (VLMs), such as LLaVA [45], GPT-4V [55], and Gemini [63], have shown impressive instruction-following abilities across various tasks [76, 43, 11, 15, 75, 33] by integrating large language models (LLMs) [65, 23] with visual encoders [14, 58]. However, VLMs may sometimes produce undesirable outputs, possibly due to omitting important details in images or misunderstanding the instructions, which prompts the need for the feedback-refining capability beyond the normal instruction-following ability. This capability enables VLMs to spontaneously refine their responses based on user feedback, as depicted in Fig. 1, enhancing the efficiency and smoothness of interactions between users and visual assistants.
@@
In this paper, we build FIRE, a dataset for Feedback Integration and Refinement Evaluation of VLMs. FIRE is composed of 1.1M high-quality multi-turn feedback-refinement conversations, derived from $27$ source datasets across a wide range of tasks, such as visual question answering [18], image captioning [8], OCR reasoning [54, 60], document understanding [20], math reasoning [47], and chart analysis [51]. To scale up the data collection, FIRE is collected in two stages. In the first stage, we randomly sample $\sim$100K image-instruction-response triplets from data sources. We use each triplet to instruct GPT-4V to simulate a dialogue between a student and a teacher: the student answers the question and the teacher provides feedback to help the student improve its answer. We filter out generated low-quality conversations, such as those with too many turns or no improvement, rendering 100K high-quality feedback-refinement conversations, named FIRE-100K. In the second stage, we fine-tune two LLaVA-NeXT [44] models on FIRE-100K: one is trained as a student to refine its answer with the feedback, and the other is trained as a teacher to generate feedback for the student’s answer. We simulate dialogues between the student and the teacher models using $\sim$1M data points from the data sources, rending a split named FIRE-1M. In this case, the full FIRE dataset consists of 1.1M feedback-refinement conversations in two splits FIRE-100K and FIRE-1M.
@@
To comprehensively evaluate the feedback-refining capability of VLMs, we build FIRE-Bench that has 11K feedback-refinement conversations derived from 16 source datasets, including test splits from 8 seen datasets in FIRE-100K and FIRE-1M, as well as 8 unseen datasets from recently-proposed popular multimodal benchmarks. Using FIRE-Bench, we design two evaluation settings: fixed dialogues and free dialogues. In fixed dialogues, we compare the model’s refined response with ground truth in the generated conversations in FIRE-Bench, given a fixed dialogue history. In free dialogues, we let the model freely interact with a teacher model about instructions in FIRE-Bench, and test how fast & how much the model can improve its answers based on the feedback provided by the teacher model.
@@
We develop FIRE-LLaVA by fine-tuning LLaVA-NeXT [44] on FIRE-100K and FIRE-1M. The evaluation results on FIRE-Bench shows that FIRE-LLaVA exhibits significant improvements based on feedback in conversations, exceeding the original LLaVA-NeXT model by $50\%$. These results underscore the significance of FIRE-100K and FIRE-1M in enhancing feedback integration, while FIRE-Bench provides an evaluation platform to analyze refinements. We expect that FIRE could motivate future exploration of the feedback-refining capability of VLMs.
@@
In summary, our contributions are three-fold. (1) We introduce FIRE, a dataset containing 1.1M feedback-refinement conversations across a wide range of tasks, where 100K data is generated by GPT-4V and 1M data is freely generated by simulating dialogues between tuned open-source models. (2) We introduce the FIRE-Bench benchmark, composed of 11K conversations and a teacher model, providing comprehensive evaluations for the feedback-refining capability in two settings: fixed dialogues and free dialogues. (3) We develop FIRE-LLaVA, an advanced VLM that could improve its responses based on feedback, making efficient interaction between users and VLMs.
@@ 2.1 Vision Language Models
Building open-source VLMs to compete with closed-source models like GPT-4V [55] and Gemini [63] is a hot research topic. BLIP [36, 35] and Flamingo [1] are pioneering models that combine LLMs with visual encoders to enhance cross-modal understanding and reasoning abilities. LLaVA [45], InstructBLIP [13], MMICL [73], and MiniGPT4 [76] develop the instruction tuning ability of VLMs by introducing a large number of instruction-response pairs. Along this way, some work focuses on the visual grounding or editing ability of VLMs [7], such as Kosmos-2 [57], SearchVLMs [34], MINI-GPTv2 [6], Qwen-VL [3], and UltraEdit [74], improving the region understanding for VLMs. InternVL [11] and mini-Gemini [38] develop powerful visual encoders for high-resolution images, and CuMo adopts a mixture-of-experts (MOE) architecture to manage diverse data better. Compared with existing VLMs, our FIRE-LLaVA has a more powerful feedback-refining capability across diverse tasks, which can spontaneously refine responses based on user feedback, leading to efficient and smooth interaction with users.
@@ 2.3 Feedback Learning in Multimodal Models
Learning from feedback is a promising research direction, playing an important role in human-robot interaction [39, 12]. Existing feedback learning methods can be roughly divided into two categories: planned feedback learning and impromptu feedback learning. Planned feedback learning updates models based on user feedback, and thus can generalize to new data but cannot provide refined responses immediately. CLOVA [17] and Clarify [30] are representative methods that automatically collect data to learn new knowledge. LLaVA-RLHF [62] collects human preference and trains VLMs via reinforcement learning. Self-refine[50] shows that LLMs could improve their responses by iteratively refining their outputs based on self-generated feedback. Impromptu feedback learning can immediately refine responses but have less generalization since they usually do not update models, which is widely studied in LLMs [2, 37, 64]. Liao et al. [40] use VLMs themselves as verifiers that produce feedback to correct recognition results. VolCaNo [29] generates data specifically for refinement to address visual hallucinations. DRESS [10] generates helpfulness, honesty, and harmlessness responses via impromptu feedback learning. Different from DRESS, we improve the correctness and details of responses via impromptu feedback learning across diverse tasks.
@@ 3.1 Task Definition
Data Source. To enhance the diversity and comprehensiveness of our dataset, we compile more than 1.1M image-instruction-response triples from $27$ source datasets (more details can be found in Appendix B), being used to generate FIRE-100K, FIRE-1M, and FIRE-Bench, as shown in Fig. 2. These datasets cover tasks including visual question answering, image captioning, complex reasoning, OCR, chart/table/document analysis, math problems, science question answering etc.
@@ 4.2 Teacher Model
LLaVA-NeXT-8B 65.9 79.0 52.0 69.8 77.3 78.5 74.4 1546.0 331.4 44.9
FIRE-LLaVA 65.8 82.9 59.8 68.4 76.8 81.5 78.5 1534.8 321.1 45.3
@@ 5.5 Ablation Studies
In Fig. 7, we present the performance curve in FIRE-Bench concerning the number of turns in dialogues, evaluating LLaVA-NeXT, FIRE100K-LLaVA, and FIRE-LLaVA. We report the percentage of correctly answered samples (those scores greater than 8) after each turn. As the number of turns increases, the percentage of correctly answered samples rises across all three models. $46.57\%$ and $46.77\%$ of the test data is correctly answered in the first turn, for the LLaVA-NeXT model and FIRE100K-LLaVA respectively. For FIRE-LLaVA, $49.60\%$ of the data is correctly answered in the first turn, and this increases to $69.19\%$ after five turns, with $19.59\%$ of the samples being corrected based on feedback. Compared to the LLaVA-NeXT model, FIRE-LLaVA shows an additional $6.79\%$ improvement (from $49.60\%\to 69.19\%$ vs. $46.57\%\to 59.37\%$), highlighting the effectiveness of FIRE-LLaVA when trained on FIRE.
@@ 6 Conclusion
Acknowledgements. This work was partly supported by the National Science and Technology Major Project (2022ZD0114900). This work was partly supported by the Natural Science Foundation of China (NSFC) under Grants No. 62176021 and No. 62172041, the Natural Science Foundation of Shenzhen under Grant No. JCYJ20230807142703006, and the Key Research Platforms and Projects of the Guangdong Provincial Department of Education under Grant No.2023ZDZX1034. Mehrtash Harandi is supported by funding from the Australian Research Council Discovery Program DP230101176.
@@ A.1 Human verification on data quality
Figure A1: The screenshot of the interface for the human verification on data quality.
A.2 Human verification on FIRE-LLaVA
To evaluate the models qualitatively, we conducted a human study comparing responses from LLaVA-Next-8B and FIRE-LLaVA. The interface is shown in Fig. A2. We randomly sampled 100 instances and provided each model with identical initial responses and feedback, asking them to generate refined responses. Three independent human evaluators assessed these responses, without knowing which model generated which response (responses were shuffled to ensure blinding). The evaluation results, detailed in Tab. A2, show that FIRE-LLaVA outperforms LLaVA-Next-8B with a significantly higher preference score (37.67 vs. 24.33), indicating that FIRE-LLaVA’s responses are more aligned with human preferences.
Table A2: Human evaluation for LLaVA-Next-8B and FIRE-LLaVA.
FIRE-LLaVA is Better Tie LLaVA-Next-8B is Better
Votes 37.67 38 24.33
Figure A2: The interface of human evaluation on FIRE LLaVA. We randomly sampled 100 instances, allowing both the baseline and FIRE-LLaVA models to generate inferences under identical former responses and feedback conditions. Three participants were invited to rate the responses. For each sample, the responses were shuffled to randomize the association between the responses and Model A/B. Participants selected their preference by choosing ‘A is better,’ ‘Tie,’ or ‘B is better’ based on their judgment.
Appendix B Data source
Our dataset uses images from 27 diverse sources to provide a robust training dataset for FIRE. All 27 datasets are public datasets, and all the images can be downloaded via links in Tab. A3. The comprehensive list of the source datasets and links to their metadata are detailed below:
Table A3: Data utilized from 27 source datasets for training and test data in FIRE.
Appendix C Additional Experimental Results
C.1 Error bar
@@
C.2 More VLMs
C.2.1 LLaVA-Next-Vicuna-7B
@@
C.2.2 LLaVA-1.5-Vicuna-7B
We have also performed experiments using LLaVA-1.5-Vicuna-7B as another baseline. The results are presented in Tab. A6. The findings demonstrate that the LLaVA-1.5-Vicuna-7B model fine-tuned on FIRE100K outperforms the original LLaVA-1.5-Vicuna-7B model across all four metrics—AT, ADR, ATR, and RR—highlighting the effectiveness of our FIRE dataset.
Table A6: Results of LLaVA-1.5-Vicuna-7B and LLaVA-1.5-Vicuna-7B-FIRE on FIRE-Bench.
Model AT ($\downarrow$) ADR ($\uparrow$) ATR ($\uparrow$) RR ($\uparrow$)
LLaVA-1.5-Vicuna 1.00 0.62 0.46 0.12
FIRE-LLaVA-1.5 0.94 0.80 0.61 0.20
C.3 More Visualization
We provide more cases tested on the FIRE-Bench to show the feedback-refining ability, as shown in Figs. A3, A4, A5, A6, A7, A8, A9 and A10. We observe that FIRE-LLaVA trained on FIRE has good feedback-refining ability on different tasks, such as visual perception, visual question answering, image captioning, OCR reasoning, document understanding, math reasoning, science question answering, and chart analysis.
@@ C.3 More Visualization
Appendix D Prompt
D.1 Prompt for GPT-4V
@@
D.2 Prompt for Student and Teacher models
@@
Appendix E Potential Negative Societal Impacts